OpenStudy (anonymous):

explain how you decide which inverse operation to use first when solving a two step equation

OpenStudy (experimentx):

for eg??

OpenStudy (anonymous):

what

OpenStudy (experimentx):

example?? what kind of equations?

OpenStudy (anonymous):

-1+2x=7

OpenStudy (anonymous):

You did not answer what is a two step equation?

OpenStudy (experimentx):

well, ... first you remove additive terms i.e including + and - 's and you remove multiplicative terms i.e. x and / here you remove -1 first so you make 2x alone =>> 2x = 7+1 then remove 2 to make x alone => x = 8/2 = 4